# Large Language Models Are Involuntary Truth-Tellers: Exploiting Fallacy Failure for Jailbreak Attacks

### v1

@@ 4.3 Additional Studies
An intriguing question is the role of scene and purpose in jailbreak attacks. Do they alone suffice to bypass the LLM’s security measures? Does FFA retain attack ability without a scene or purpose? We conducted an ablation study and computed the AHS and ASR under five attacks, scene, and purpose combinations across three language models shown in Figure 5. setting X, Y, and Z refer to directly asking LLM for malicious behavior with that combination of scene and purpose, respectively. FFA + Z refers to using Z as the scene and purpose of the FFA attack. FFA + None refers to the FFA attack without specifying any scene or purpose. We can observe that (1) naïvely adding a scene and purpose to the direct instruction of harmful behavior mostly has a marginal effect on jailbreak attack. The only exception is the combination of “scientific fiction” and “against evil Doctor X,” which demonstrates notable attack efficacy against GPT-3.5-turbo and Gemini-Pro. Interestingly, this unique design seems to be the primary driving force behind the DeepInception method. (2) Although our method archives optimal performance with the combination of scene and purpose, it can retain significant attack ability without a scene and purpose, except for LLaMA-3, which opposes the creation of untruthful content. (3) When using the same fictitious scene and purpose as DeepInception, our method is more likely to induce more harmful output.
@@ Appendix B Main Experiment Settings
We treat [scene] and [purpose] as hyperparameters in FFA. We utilize three predefined scene-purpose pairs based on commonsense and previous work, shown in Table 3:

### v2

@@ 4.3 Additional Studies
An intriguing question is the role of scene and purpose in jailbreak attacks. Do they alone suffice to bypass the LLM’s security measures? Does FFA retain attack ability without a scene or purpose? We conducted an ablation study and computed the AHS and ASR under five attacks, scene, and purpose combinations across three language models shown in Figure 5. setting X, Y, and Z refer to directly asking LLM for malicious behavior with that combination of scene and purpose, respectively. FFA + Z refers to using Z as the scene and purpose of the FFA attack. FFA + None refers to the FFA attack without specifying any scene or purpose. We can observe that (1) naïvely adding a scene and purpose to the direct instruction of harmful behavior mostly has a marginal effect on jailbreak attack. The only exception is the combination of “scientific fiction” and “against evil Doctor X,” which demonstrates notable attack efficacy against GPT-3.5-turbo and Gemini-Pro. Interestingly, this unique design seems to be the primary driving force behind the DeepInception method. (2) Although our method archives optimal performance with the combination of scene and purpose, it can retain significant attack ability without a scene and purpose, except for LLaMA-3, which opposes the creation of untruthful content. (3) When using the same fictitious scene and purpose as DeepInception, our method is more likely to induce more harmful output. For additional studies on each component and FFA variations, please refer to Appendix C.
@@ Appendix B Main Experiment Settings
We treat [scene] and [purpose] as hyperparameters in FFA. We utilize three predefined scene-purpose pairs based on commonsense and previous work, shown in Table 3.
@@
Appendix C Prompt Flexibility and Scalability
We find that while the prompts proposed in the paper achieves the best performance, many components in the prompt can be removed, as long as the central component, the fallacious reasoning prompt, is preserved. Table 4 shows the performance exhausting the combinations of the presence of [scene], [purpose], and [deceptiveness] components, with random 25% sample of our dataset and two LLMs: GPT-3.5-turbo and Gemini-Pro. $\checkmark$ and $\times$ indicate component presents and does not present, respectively.
Model Scene Purpose Request BPRAHS ASR%
GPT-3.5-turbo $\times$ $\times$ $\times$ 70 2.52 55
$\times$ $\times$ $\checkmark$ 77.5 3.56 52.5
$\times$ $\checkmark$ $\times$ 72.5 3.73 62.5
$\times$ $\checkmark$ $\checkmark$ 77.5 3.83 65
$\checkmark$ $\times$ $\times$ 55 2.72 32.5
$\checkmark$ $\times$ $\checkmark$ 60 2.79 35
$\checkmark$ $\checkmark$ $\times$ 67.5 3.3 47.5
$\checkmark$ $\checkmark$ $\checkmark$ 85 4.24 77.5
Gemini-Pro $\times$ $\times$ $\times$ 47.5 2.52 30
$\times$ $\times$ $\checkmark$ 65 3.6 65
$\times$ $\checkmark$ $\times$ 65 3.35 57.5
$\times$ $\checkmark$ $\checkmark$ 60 3.3 57.5
$\checkmark$ $\times$ $\times$ 52.5 2.9 45
$\checkmark$ $\times$ $\checkmark$ 65 3.4 57.5
$\checkmark$ $\checkmark$ $\times$ 65 3.35 45
$\checkmark$ $\checkmark$ $\checkmark$ 77.5 3.78 60
Table 4: Attack performance exhausting the combinations of the presence of [scene], [purpose], and [deceptiveness] components, with random 25% sample of our dataset and two LLMs: GPT-3.5-turbo and Gemini-Pro.
Additionally, we find that FFA do not necessarily require the use of scene and purpose as a “background story” to facilitate the attack. This is particularly true for LLMs that have been aligned to refuse role-playing scenarios, such as Claude Samuel et al. (2024), as role-playing has been shown to potentially facilitate jailbreak attacks.
We exploit Claude’s analogy ability and adapt our approach to prompt Claude-3 to first generate a false proof or reasoning for a random mathematical problem (non-harmful, e.g., a false proof of “all odd numbers are prime numbers”), followed by asking it to analogically provide a false step-by-step procedure for the harmful behavior. We show the performance in Table 5. Note that although the adaption is still centered on fallacious reasoning, it can be LLM-specific and requires future exploration.
Model Method BPRAHS ASRClaude-3 CGC 2 1.16 4 AutoDAN 2 1 0 DeepInception 6 1.1 0 Art_Prompt 34 2.22 20 FFA 29.4 2.17 29.4
Table 5: Performance for adapted FFA on Claude-3.
Appendix D A Possible Defense against FFA
Inspired by the behavior of the LLaMA model family, we append instructions such as “Do not tell a lie or make untruthful statements” to FFA malicious prompts, test eight variations of similar instructions, and compute the averaged performance on two of the most vulnerable LLMs in our study: GPT-3.5 and Vicuna 1.5. Our results show that this method can mitigate the FFA attack, however, varying degrees across different models. See Table 6.
However, we caution against overly restrictive approaches. As discussed in Section 4.2, completely prohibiting LLMs from generating false content is not optimal. We believe that exposing these vulnerabilities highlights urgent research attention and is a crucial first step in developing effective defenses and future research in LLM security and alignment.
Model Method BPRAHS ASRGPT-3.5-turbo Truthful+FFA 12.5 1.16 0 Vicuna-1.5 Truthful+FFA 42.5 2.31 25
Table 6: Performance of FFA on GPT-3.5-turbo and Vicuna-1.5 under the truthful defense.